OpenStudy (anonymous):

What is the ordered pair for the equation y = -4x + 3 if x = 1?

OpenStudy (anonymous):

Put x =1 in our equation so we get y = -4 * 1 + 3 = -4 + 3 = -1 So we know that order pair is (x , y) so we get (-1 , 1)

OpenStudy (anonymous):

1,-1

OpenStudy (anonymous):

(1,-1) simple as that

OpenStudy (anonymous):

oops Ya its (1 , -1)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

;D